# Off-Path TCP Hijacking in Wi-Fi Networks: A Packet-Size Side Channel Attack

### v1

@@ 6.2 Analysis of AP Routers
The early WEP used RC4 and MIC algorithms for data encryption, whereas WPA2 replaced them with the AES-CCMP algorithm. In the latest WPA3 standard, AES-GCMP is proposed to be used as the encryption method for WPA3 Enterprise mode. However, none of these security mechanisms can prevent the encrypted frame size from leaking upper layer information. Based on our empirical findings, we confirm that all 30 evaluated wireless routers could not protect the supplicant from our attacks.
@@
Due to space limitations, Table 4 presents a concise overview of routers from the 9 AP vendors included in our investigation. Comprehensive information on all tested routers can be found in Appendix A (see Table 5). Take the first line for example, the evaluated router “Mi 4C” manufactured by Xiaomi belongs to the older Wi-Fi generation (Wi-Fi 4) and lacks support for IPv6 and WPA3. As outlined in the product description, the “Mi 4C” device offers support for various security features. These include a built-in firewall that allows administrators to define packet forwarding rules, flood defense mechanisms that restrict malicious flood traffic to prevent DoS attacks, and MAC address filtering, which enables network access authorization based on hardware addresses. In our investigation, all vulnerable routers claim to support different security mechanisms to prevent various attacks. However, our study demonstrates that the existing security mechanisms are inadequate against our attack.
@@ 7 Discussion
7.1 Success Rate Analysis
Our attack relies on the observation of the victim’s encrypted frame size. However, the attacker’s ability to monitor the victim’s frames may be hindered by wireless interference and Wi-Fi channel contention. These factors can directly affect the success and effectiveness of our attack. Furthermore, background traffic on the victim’s device and shifting of the TCP receive window can also influence the attack’s success.
@@ 7.1 Success Rate Analysis
Channel Contention. APs and supplicants based on the 802.11 standards use Carrier Sense Multiple Access with Collision Avoidance (CSMA/CA) to compete equally for the occupation of the wireless channel. Before transmitting frames, wireless channel listening is conducted to ensure that the channel is not occupied. Frames are transmitted only after verifying the channel’s availability. Due to channel contention, there is an uncertain delay in the victim’s responses to the probe packets. This uncertain response delay is the main reason for the fluctuating success rate of our attack because the attacker needs to analyze the victim’s encrypted frames within a time slice after the probe packets are sent. To eliminate this uncertain response delay, we propose a channel “eviction” strategy. The attacker can evict other supplicants from the channel used by the victim. Specifically, the attacker impersonates the AP and sends decertification frames to the supplicant, causing it to detach from the current channel of the AP. The supplicant will then request to access the Wi-Fi network again, but it will switch to another channel. This strategy requires the Wi-Fi network to support multiple access channels. Fortunately, most Wi-Fi networks provide more than one access channel, as shown in Sec 6.1. Note that the channel switching (i.e., our “eviction” strategy to cause other supplicants detaching from the current channel) is transparent to the users. The only impact is that the user may experience a brief (a few seconds) network jitter during the channel switching.
Refer to caption (a) 40 pkt/s of empty ACK traffic at various time intervals. Refer to caption (b) Empty ACK traffic with different send rates at 8s time intervals.
Figure 10: Quantitative analysis of empty ACK traffic.
Background Traffic. The background traffic may degrade the quality of the side channel (i.e., victim’s frame size), thereby impacting the effectiveness of the attack. Specifically, other packets received by the victim client have the potential to affect our side channel attack in two ways. (i) The Background traffic may overshadow the target encrypted frames. During our evaluation, the number of responses triggered by the probe packets is typically less than 10% compared to the background traffic. To effectively filter the target frames from the background traffic, the attacker can employ multiple WNICs to monitor the Wi-Fi channel. (ii) The responses triggered by the probe packets may be mixed with other TCP packets. If the TCP packets in the background traffic have the same size (e.g., empty ACK packets) as the response from the victim server, the attacker may mistake them for the actual responses. To quantitatively evaluate the impact of empty ACK packets on our attack, we send empty ACK packets (from other servers) with varying bandwidths (e.g., different time intervals and send rates) to the victim’s device. Simultaneously, the attacker infers the 4-tuple, sequence number, and acknowledgment number of the victim’s TCP connection by following the attack steps delineated in Section 4. The attacker repeats the attack 15 times under each experimental configuration. Inferring the correct TCP connection information is considered a successful attack. The experimental results are depicted in Figure 10. As depicted in Figure 10(a), when the time interval between empty ACK traffic decreases, the attack requires more time to execute and yields a reduced success rate. This phenomenon occurs because the presence of empty ACK packets confuses the attacker, leading to repeated verification attempts to confirm the inferred value’s accuracy. When the send rate of empty ACK packets increases (as shown in Figure 10(b)), the time required for the attack to succeed extends and the success rate diminishes, but not significantly. This phenomenon arises because our attack relies on the presence or absence of frames of specific dimensions, rather than their quantity. Consequently, the attack is more susceptible to the temporal intervals at which empty ACK packets are observed. Fortunately, in most cases (e.g., content servers), the TCP packets sent by the servers contain application data. As a result, TCP packets from background traffic are significantly larger (over 1000 bytes) in size compared to the responses (e.g., challenge ACK) triggered by the probe packets. Additionally, in scenarios with a high volume of empty ACK traffic affecting the victim, the attacker has the option to leverage SACK-ACK to infer the port number and acknowledgment number of the TCP connection, thus avoiding interference from empty ACKs. Specifically, (i) when inferring the port number, the attacker sends two TCP packets containing data to the server, each bearing sequence numbers $seq$ and $seq+2^{31}$ respectively. If the TCP port is accurately inferred, the attacker will encounter an 80-byte encrypted frame, as one of the two packets in question is bound to elicit the server’s SACK-ACK response. Conversely, if the inference is incorrect, the attacker will not observe the 80-byte encrypted frame. (ii) Since TCP is full duplex, the attacker can utilize SACK-ACK to infer the sequence number on the client side and consequently obtain the acknowledgment number on the server side.
Shifting Receive Window. An additional factor influencing the attack’s success rate is the dynamic adjustment of the receive window. When the victim’s TCP connection carries on ongoing traffic, the acceptable sequence and acknowledgment windows will shift during the attack, but it does not impact the attack as long as the inferred sequence number and acknowledgment numbers fall in the sliding windows. The attacker can repeatedly infer the sequence number and acknowledgment number. Even if the receive window slides quickly enough to thwart the attacker’s inference, the attacker can opt to target the other end of the TCP connection. In this scenario, the receive window adjusts at a slower pace, as has been demonstrated in prior research [21, 14].
7.2 Special Policies Analysis
In practice, two special network policies (i.e., frame aggregation and AP isolation) may also affect our attack.
@@ 7.2 Special Policies Analysis
Frame Aggregation. The MAC layer frame aggregation technique is proposed in the 802.11n [2] standard to improve the throughput and efficiency of WLANs. There are two methods available to perform frame aggregation, i.e., aggregate MAC protocol service unit (A-MSDU) and aggregate MAC protocol data unit (A-MPDU). The main difference between MSDU and MPDU is that the latter has a MAC header through 802.11 protocol encapsulation while the former becomes MPDU after adding integrity check MIC, encryption, sequence number assignment, CRC checksum, and MAC header. The A-MPDU has no impact on our attack because each MPDU has a complete MAC header and the attacker can distinguish the encryption payload size of each MPDU. However, if the victim triggers A-MSDU, multiple packets will be encrypted together, preventing the attacker from inferring TCP information based on the encrypted frame size. Fortunately, the server’s responses triggered by attackers are rarely aggregated into A-MSDUs. In the following, we analyze the reasons why A-MSDU frames are not triggered.
@@
The A-MSDU completes when the size of the waiting packets reaches the maximum A-MSDU threshold or the maximum delay of the oldest packets reaches a pre-assigned value. Its maximum size can be 3839 or 7935 bytes, depending on the throughput capacity of the station (STA). The size can be found from the High-Throughput (HT) capabilities element of the HT STA release. In case the aggregated frame size does not reach the aggregation threshold, the MSDU buffer queue waits for new MSDUs to reach the MAC layer. But if the maximum delay exceeds the preset maximum, the aggregated frame will be immediately inserted into the channel, even if the aggregated frame size does not reach the aggregation threshold. The maximum delay is typically set to 1 µs  [55]. Additionally, only frames with the same receiver and the same TID (traffic identifier) can be aggregated together using A-MSDU. In our attack, few A-MSDU frames are observed. We speculate that this absence may be due to the attacker sending probe packets at a low rate, and the response packets having a different TID than the background traffic (e.g., video packets). Consequently, the TCP responses triggered by the probe packets are not aggregated into A-MSDU.
AP Isolation. When AP isolation is enabled on wireless routers, supplicants are only allowed to access the gateway and the outside network. This means that supplicants in the same Wi-Fi network cannot communicate with each other. Under this circumstance, the probe packets sent by the attacker to the victim supplicant will be discarded at the AP. This particular network policy can throttle our attack. However, AP isolation may impair network availability and impact user experience. For example, the wireless display can no longer be used in Wi-Fi networks when AP isolation is enabled. As a result, we observe that less than 14% (11 out of 80) of the evaluated Wi-Fi networks enforce this policy.
@@ 8 Countermeasure
Defenses in 802.11 Standard. As Wi-Fi networks rely on shared wireless media, any 802.11-compliant device has the capability to sniff all Wi-Fi frames. To maintain the confidentiality and integrity of Wi-Fi frames, encryption mechanisms are commonly employed in Wi-Fi networks. Although Wi-Fi networks encrypt their frames transmitted in the wireless channel, there exists a strong correlation between the encrypted frame size and the upper layer applications. This correlation allows an off path attacker to analyze the encrypted frame size, infer the victim’s TCP information, and subsequently conduct the TCP hijacking attack. Adjusting the security mechanisms of the 802.11 standards so that the AP or supplicant dynamically pads the size of encrypted frames is one possible countermeasure. This countermeasure may require changes and redesign at the Wi-Fi standard level. We are currently in discussions with the Wi-Fi Alliance regarding this countermeasure.
@@ 9 Related Work
Wi-Fi Attacks. While Wi-Fi serves as a widely used access method for end-users to connect to the internet, it presents higher security risks compared to wired LANs, such as Ethernet. Public Wi-Fi networks, in particular, are susceptible to attacks due to their open-access nature. To safeguard wireless users in Wi-Fi networks, numerous security mechanisms have been proposed in recent years, including WEP, WPA, WPA2, and WPA3 [8]. Nevertheless, existing researches [57, 24, 39, 61, 62, 63, 60] have revealed implementation vulnerabilities or design flaws in these security mechanisms that can compromise Wi-Fi networks. For example, WPA is vulnerable to key recovery attacks [57, 39] and dictionary attacks [37]. Subsequently, WPA2 and WPA3 were introduced to mitigate these vulnerabilities. However, recent research indicates that WPA2 is susceptible to KRACK attacks [61, 62], and WPA3 can be compromised by downgrade or dictionary attacks [63]. Besides, recent studies [60, 48] have revealed that attackers can leverage the design flaws of Wi-Fi networks to circumvent these security mechanisms. Unlike the aforementioned studies, our attack does not require cracking or circumventing security mechanisms.
In addition to Wi-Fi network cracking, extensive research has been conducted on traffic hijacking within Wi-Fi networks. Attackers can execute an Evil Twins attack by deploying a rogue AP to hijack the traffic of victim supplicants [25, 28, 9, 40]. Additionally, rogue DHCP and ARP poisoning are recognized as common threats in Wi-Fi networks. Notably, these attacks have been subject to extensive research, leading to the development of countermeasures, including rogue AP and rogue DHCP detection [36, 32, 5, 10] and ARP protection [17, 50, 4]. Recently, Feng et al. exposed implementation flaws in wireless routers that allow the hijacking of traffic from victim supplicants in Wi-Fi networks using ICMP redirect messages. In contrast to previous studies, our attack is novel and stealthy by revealings a fundamental security vulnerability in the 802.11 protocol that affects all Wi-Fi networks.
@@
In TCP connection hijacking attacks, side channels serve as potent tools for attackers. The IPID, in particular, has been a frequent target for exploitation. For example, Jeffrey et al. utilized per-destination IPID counters to estimate the number of packets transmitted between two machines and even detect the presence of a TCP connection [34]. Similarly, Alexander et al. used the IPID of triggered RST packets to identify the existence of a victim TCP connection [7]. In a recent instance, Feng et al. manipulated the IPID assignment using ICMP to hijack the victim’s TCP connection [21]. Besides IPID, the challenge ACK mechanism is another side channel exploited by off-path attackers. Cao et al., for example, utilized the global rate limit of challenge ACK to infer and hijack TCP connections [14, 15]. Moreover, a timing side channel has been found in half-duplex Wi-Fi technology, which can be exploited by off-path attackers to inject data into a TCP connection [16]. Tolley et al. recently proposed a blind in/on-path attack in VPNs, aiming to infer the existence of, interfere with, or inject data into TCP connections forwarded through encrypted VPN tunnels [58]. Different from previous research, our work reveals a new side channel, i.e., information leakage due to the encrypted frame size in Wi-Fi networks. This side channel can be exploited by a pure off-path attacker to hijack victim’s TCP connections.

### v4

@@ 6.1 Analysis of AP Routers
The early WEP used RC4 algorithm for data encryption, whereas WPA2 replaced them with the AES-CCMP algorithm. In the latest WPA3 standard, AES-GCMP is proposed to be used as the encryption method for WPA3 Enterprise mode. However, none of these security mechanisms can prevent the encrypted frame size from leaking upper layer information. Based on our empirical findings, we confirm that all 30 evaluated wireless routers could not protect the supplicant from our attacks.
@@
Table 3 shows detailed information on 30 tested wireless routers in our investigation. Take the first line for example, the evaluated router “Mi 4C” manufactured by Xiaomi belongs to the older Wi-Fi generation (Wi-Fi 4) and lacks support for IPv6 and WPA3. As outlined in the product description, the “Mi 4C” device offers support for various security features. These include a built-in firewall that allows administrators to define packet forwarding rules, a flood defense mechanism that restricts malicious flood traffic to prevent DoS attacks, and MAC address filtering, which enables network access authorization based on hardware addresses. In our investigation, all tested routers claim to support different security mechanisms to prevent various attacks. However, our study demonstrates that the existing security mechanisms are inadequate against our attack.
Table 4: Experimental results in 30 real-world Wi-Fi networks.
No.
SSID
AP Vendor
IPv4/IPv6
PHY model
AP isolation
W
i-Fi channel
SSH DoS
Web hijack
1 Bookstore 1 ADSLR 802.11n/ac No 6, 161 7/10 6/10
2 Bookstore 2 HUAWEI 802.11n/ac/ax No 11, 44 7/10 7/10
3 Bookstore 3 Xiaomi 802.11n/ac No 6, 149 8/10 7/10
4 Coffee Shop 1 TP-LINK 802.11n/ac No 6, 60 8/10 6/10
5 Coffee Shop 2 Wimaster 802.11n/ac Yes 1, 48 7/10 6/10
6 Coffee Shop 3 Tenda 802.11n/ac No 4, 153 6/10 5/10
7 Restaurant 1 D-Link 802.11n/ac No 5, 149 7/10 5/10
8 Restaurant 2 Ruijie 802.11n/ac Yes 11, 64 6/10 4/10
9 Restaurant 3 iKuai 802.11n/ac No 1, 48 5/10 3/10
10 Office building 1 TP-LINK 802.11n/ac No 11, 36, 40 7/10 6/10
11 Office building 2 H3C 802.11n/ac No 1, 48, 153 8/10 7/10
12 Office building 3 Netcore 802.11n/ac Yes 6, 149 8/10 6/10
13 Enterprise 1 TP-LINK 802.11n/ac No 6, 36 6/10 6/10
14 Enterprise 2 HUAWEI 802.11n/ac Yes 11, 157 7/10 6/10
15 Enterprise 3 Ruijie 802.11n/ac Yes 1, 11, 40, 149 6/10 5/10
16 Fast Food Restaurant 1 Wimaster 802.11n/ac/ax No 6, 161, 149 6/10 4/10
17 Fast Food Restaurant 2 TP-LINK 802.11n/ac No 3, 157 7/10 6/10
18 Fast Food Restaurant 3 Ruijie 802.11n/ac No 1, 44 6/10 6/10
19 Cinema 1 HUAWEI 802.11n/ac No 1, 157 7/10 6/10
20 Cinema 2 Ruijie 802.11n No 6 7/10 6/10
21 Cinema 3 H3C 802.11n/ac No 10, 149 7/10 5/10
22 Hotel 1 HUAWEI 802.11n/ac No 6, 44 8/10 7/10
23 Hotel 2 D-Link 802.11n/ac No 1, 48 6/10 5/10
24 Hotel 3 Xiaomi 802.11n Yes 1 5/10 4/10
25 Experience Store 1 HUAWEI 802.11n/ac No 1, 36 7/10 6/10
26 Experience Store 2 HUAWEI 802.11n/ac No 11, 149 7/10 6/10
27 Experience Store 3 Tenda 802.11n/ac No 4,153 6/10 5/10
28 Campus 1 Xiaomi 802.11n/ac No 9, 36 6/10 4/10
29 Campus 2 Ruijie 802.11n/ac No 1, 44 7/10 6/10
30 Campus 3 H3C 802.11n/ac No 1, 6, 40, 64 6/10 6/10
  • ◐ means IPv4 only and ● means both IPv4 and IPv6 are supported.
6.2 Real-world Wi-Fi Networks Evaluation
The Wi-Fi scenarios we tested cover a wide range of public settings, including coffee shops, restaurants, hotels, cinemas, and bookstores. The experimental results illustrate that over 93% (i.e., 75 out of 80) of the evaluated Wi-Fi networks are vulnerable to our attack. Next, we elaborate on the evaluation results.
As shown in Figure 10, out of the 80 Wi-Fi networks we assessed, 74 are found to be IPv4-only networks, while the remaining 6 have IPv6 capabilities [^8] Despite the accelerated deployment of IPv6 networks in recent years [6], the adoption of IPv6 support in Wi-Fi networks remains relatively limited.. This can be attributed, in part, to the lack of IPv6 support in legacy wireless routers and the limited incentive for merchants to invest in new wireless routers.
Figure 10: Attack evaluation on 80 real-world Wi-Fi networks.
The 802.11n/ac standards are predominantly utilized (73.75%) in real-world Wi-Fi networks. This indicates that these Wi-Fi networks support two frequency bands (i.e., 2.4 GHz and 5 GHz), with their physical layer models based on the 802.11n and 802.11ac standards, respectively. There is only 17.5% (14 out of 80) of the evaluated Wi-Fi networks support 802.11ax. This is consistent with our expectations, as the 802.11ax standard was defined in 2019 and would require more time for widespread deployment. Furthermore, the simultaneous utilization of two channels in Wi-Fi networks is the most prevalent case (76.25%) due to the widespread support and default configuration of dual-channel capabilities in wireless routers. In certain scenarios, such as office buildings, Wi-Fi networks employ multiple wireless channels to enhance network performance. In our study, we identify 12 Wi-Fi networks that utilize multiple wireless channels. At first glance, the usage of multiple wireless channels might appear as a minor hurdle to our attack, as the attacker needs to perform additional channel scanning to determine the specific channel employed by the victim. Indeed, the attacker can enhance the success rate by employing a channel “eviction” strategy, which will be explained in detail in Section 7. Additionally, we encounter several Wi-Fi networks that operate on a single channel. When questioned, network administrators cited security considerations as the rationale behind this choice, although they did not provide any further specifics.
Out of all the evaluated networks, 16 (20%) of them are open and do not encrypt the supplicant’s data frames. This means that an attacker can potentially access and view the contents of all supplicant’s frames transmitted on these networks, representing a significant breach of supplicant privacy. The remaining 64 (80%) Wi-Fi networks utilize WPA2/WPA3 to encrypt the supplicant’s wireless frame. It is worth noting that out of the 80 Wi-Fi networks evaluated, 11 (13.75%) of them have AP isolation enabled. In these Wi-Fi networks, the attacker obtains the victim’s MAC and IP address by leveraging the DHCP mechanism and injects malicious packets into the victim through the AP’s external port. However, our attack encounters failure in five Wi-Fi networks. Among them, one network is equipped with reverse path authentication [51, 58], preventing the attacker from sending packets from the WLAN to the AP’s external port. In the remaining four Wi-Fi networks, the attacker cannot obtain the AP’s external port as the gateway does not respond to ICMP ping messages.
We elaborate more on the experimental results of 30 encrypted Wi-Fi networks in Table 4. We take the first row of Table 4 as an example to analyze the results. In our study, the SSID “Bookstore 1” indicates a Wi-Fi network that is accessible in a bookstore. It is common practice to set the Wi-Fi SSID as the organization name, which may expose the organization’s identity. Therefore, to protect anonymity, we have anonymized the Wi-Fi SSID in this paper. This bookstore’s Wi-Fi network only supports IPv4 and does not have AP isolation enabled, while its AP is produced by ADSLR. This AP provides two access channels (i.e., 6 and 161) and employs the 802.11n and 802.11ac standards. The TCP connection of the victim supplicant can be hijacked using the attack presented in Section 4. The success rates for conducting SSH DoS and web hijacking on the victim supplicant are 70% and 60%, respectively.
Within the evaluated vulnerable Wi-Fi networks, we have observed a range of success rates for our attack, varying from 30% to 80%. The principal factor influencing this variance is the heterogeneous wireless environments (e.g., different wireless interference and channel contention) encountered in real-world Wi-Fi networks, leading to varying capabilities for attacker to capture the victim’s Wi-Fi frames. Factors such as wireless interference (e.g., from microwave ovens and Bluetooth devices) and channel contention can hinder the attacker’s ability to capture the victim’s Wi-Fi frames, resulting in the failure of attacks. For instance, as illustrated in Table 4, Coffee Shop 1, situated on a university campus, experiences less wireless interference and channel contention compared to Restaurant 3, located within a large shopping mall. Consequently, the success rate of attacks in the latter Wi-Fi network is lower due to the elevated interference and contention in that environment. We will conduct a more comprehensive analysis of the factors influencing the success of our attack in Section 7.
@@ 7 Discussion
Our attack relies on the observation of the victim’s encrypted frame size. However, the attacker’s ability to monitor the victim’s frames may be hindered by wireless interference, Wi-Fi channel contention and frame aggregation. These factors can directly influence the success and effectiveness of our attack. We delve into the details of these factors in this section.
@@
Channel Contention. APs and supplicants based on the 802.11 standards use Carrier Sense Multiple Access with Collision Avoidance (CSMA/CA) to compete equally for the occupation of the wireless channel. Before transmitting frames, wireless channel listening is conducted to ensure that the channel is not occupied. Frames are transmitted only after verifying the channel’s availability. Due to channel contention, there may be an uncertain delay or even frame dropping in the victim’s responses to the probe packets. This uncertain response delay or frame dropping is the primary reason for the fluctuating success rate of our attack because the attacker needs to analyze the victim’s encrypted frames within a time slice after the probe packets are sent. To mitigate channel contention, we propose a channel “eviction” strategy. The attacker can evict other supplicants from the channel used by the victim. Specifically, the attacker impersonates the AP and sends decertification frames to the supplicant, causing it to detach from the current channel of the AP [^9] In the Wi-Fi network with MFP (Management Frame Protection) enabled, attackers can exploit implementation vulnerabilities to force the supplicant to detach from the current channel [49, 48].. The supplicant will attempt to reconnect to the Wi-Fi network, but after encountering several disconnections, it will switch to another channel. This strategy requires the Wi-Fi network to support multiple access channels. Fortunately, most Wi-Fi networks provide more than one access channel, as shown in Section 6.2. Note that the channel switching (i.e., our “eviction” strategy to cause other supplicants to detach from the current channel) is transparent to the users. The only impact is that the user may experience a brief (a few seconds) network jitter during the channel switching.
@@
Frame Aggregation. The MAC layer frame aggregation technique is proposed in the 802.11n standard [2] to improve the throughput and efficiency of WLANs by combining multiple data packets into a single transmission unit. There are two methods available to perform frame aggregation, i.e., aggregate MAC protocol service unit (A-MSDU) and aggregate MAC protocol data unit (A-MPDU). The main difference between MSDU and MPDU is that the latter has a MAC header through 802.11 protocol encapsulation while the former becomes MPDU after adding integrity check MIC, encryption, sequence number assignment, CRC checksum, and MAC header. The A-MPDU has no impact on our attack because each MPDU has a complete MAC header and the attacker can distinguish the encryption payload size of each MPDU. However, if the victim triggers A-MSDU, multiple packets will be encrypted together, preventing the attacker from inferring TCP information based on the encrypted frame size. Fortunately, the server’s responses triggered by attackers are rarely aggregated into A-MSDUs. In the following, we analyze the reasons why A-MSDU frames are not triggered.
@@
The A-MSDU completes when the size of the waiting packets reaches the maximum A-MSDU threshold or the maximum delay of the oldest packets reaches a pre-assigned value. Its maximum size can be 3839 or 7935 bytes, depending on the throughput capacity of the station (STA). The size can be found from the High-Throughput (HT) capabilities element of the HT STA release. In case the aggregated frame size does not reach the aggregation threshold, the MSDU buffer queue waits for new MSDUs to reach the MAC layer. But if the maximum delay exceeds the preset maximum, the aggregated frame will be immediately inserted into the channel, even if the aggregated frame size does not reach the aggregation threshold. The maximum delay is typically set to 1 µs  [57]. Additionally, only frames with the same receiver and the same Traffic Identifier (TID) can be aggregated together using A-MSDU. In our attack, few A-MSDU frames are observed. We speculate that this absence may be due to the attacker sending probe packets at a low rate (compared to 1 µs) [^10] The attacker can control the time interval between sending probe packets to be greater than 1 µs., and the response packets having a different TID than the background traffic (e.g., video packets). Consequently, the TCP responses triggered by the probe packets are not aggregated into A-MSDU.
@@ 8 Countermeasure
Defenses in 802.11 Standard. As Wi-Fi networks rely on shared wireless media, any 802.11-compliant device has the capability to sniff all Wi-Fi frames. To maintain the confidentiality and integrity of Wi-Fi frames, encryption mechanisms are commonly employed in Wi-Fi networks. Although Wi-Fi networks encrypt their frames transmitted in the wireless channel, there exists a strong correlation between the encrypted frame size and the upper layer applications. This correlation allows an off-path attacker to analyze the encrypted frame size, infer the victim’s TCP information, and subsequently conduct the TCP hijacking attack. Adjusting the security mechanisms of the 802.11 standards so that the AP or supplicant dynamically pads the size of encrypted frames is one possible countermeasure. This countermeasure may require changes and redesign at the Wi-Fi standard level. We are currently in discussions with the Wi-Fi Alliance regarding this countermeasure.
@@ 9 Related Work
Wi-Fi Attacks. While Wi-Fi serves as a widely used access method for end-users to connect to the internet, it presents higher security risks compared to wired LANs, such as Ethernet. Public Wi-Fi networks, in particular, are susceptible to attacks due to their open-access nature. To safeguard wireless users in Wi-Fi networks, numerous security mechanisms have been proposed in recent years, including WEP, WPA, WPA2, and WPA3 [8]. Nevertheless, existing researches [60, 25, 40, 64, 65, 66, 63] have revealed implementation vulnerabilities or design flaws in these security mechanisms that can compromise Wi-Fi networks. For example, WPA is vulnerable to key recovery attacks [60, 40] and dictionary attacks [38]. Subsequently, WPA2 and WPA3 were introduced to mitigate these vulnerabilities. However, recent research indicates that WPA2 is susceptible to KRACK attacks [64, 65], and WPA3 can be compromised by downgrade or dictionary attacks [66]. Besides, recent studies [63, 48] have revealed that attackers can leverage the design flaws of Wi-Fi networks to circumvent these security mechanisms. Unlike the aforementioned studies, our attack does not require cracking or circumventing these security mechanisms.
In addition to Wi-Fi network cracking, extensive research has been conducted on traffic hijacking within Wi-Fi networks. Attackers can execute an Evil Twins attack by deploying a rogue AP to hijack the traffic of victim supplicants [26, 29, 9, 41]. Additionally, rogue DHCP and ARP poisoning are recognized as common threats in Wi-Fi networks. Notably, these attacks have been subject to extensive research, leading to the development of countermeasures, including rogue AP and rogue DHCP detection [37, 33, 5, 10] and ARP protection [17, 52, 4]. Recently, Feng et al. revealed vulnerabilities in the implementation of IP source address checking in wireless routers, enabling attackers to hijack victim’s traffic in Wi-Fi networks using ICMP redirect messages [23]. Yang et al. proposed exploiting flaws in RST packet inspection implementation in wireless routers to manipulate NAT mapping states and hijack the victim’s TCP connection [69]. In contrast, our attack does not rely on such implementation flaws in wireless routers. Instead, it reveals a novel fundamental security vulnerability in the 802.11 standards, affecting all Wi-Fi networks.
@@
In TCP connection hijacking attacks, side channels serve as potent tools for attackers. The IPID, in particular, has been a frequent target for exploitation. For example, Jeffrey et al. utilized per-destination IPID counters to estimate the number of packets transmitted between two machines and even detect the presence of a TCP connection [35]. Similarly, Alexander et al. used the IPID of triggered RST packets to identify the existence of the victim’s TCP connection [7]. In a recent instance, Feng et al. manipulated the IPID assignment using ICMP to hijack the victim’s TCP connection [22]. Besides IPID, the challenge ACK mechanism is another side channel exploited by off-path attackers. Cao et al., for example, utilized the global rate limit of challenge ACK to infer and hijack TCP connections [14, 15]. Moreover, a timing side channel has been found in half-duplex Wi-Fi technology, which can be exploited by off-path attackers to inject data into the victim’s TCP connection [16]. Tolley et al. recently proposed a blind in/on-path attack in VPNs, aiming to infer the existence of, interfere with, or inject data into TCP connections forwarded through encrypted VPN tunnels [61]. Different from previous research, our work reveals a new side channel, i.e., information leakage due to the encrypted frame size in Wi-Fi networks. This side channel can be exploited by a pure off-path attacker to hijack victim’s TCP connections.